Wilmington comapny has two manfacturing departments- Assembly and fabricarion. It considers all of its manufacturing overhead costs to be fixed costs. The first set of data thar is shown below is based on estimates from the beginning of the year. The second set of data relates to one particular job completed during the year- job bravo.

Esimated Assembly Farbrication Total

Manufacturing overhead costs $ 7,250,000 $ 7,830,000 $ 15, 080,000

Direct labor-hours 145,000   87,000 232,000

Machine - Hours 58,000 290,000 348,000

Job Bravo 30 22 52

Direct labor-hours 22 25 47

If wilmington used a plantwide predetemined overhead rate based on direct labor -hours, how much manufacturing overhead would be applied to job bravo?

If wilmington uses departmental predetemined overhead rates with direct labor-hours as the allocation bas in assembly and machine- hours as the allocation base in fabrication, how much manufacuring overhead would be applied to job bravo?    
Answer #1

a) Plantwide overhead rate = 15080000/232000 = 65 per labor hour

Overhead applied to Job Bravo = 47*65 = 3055

b) Assembly overhead rate = 7250000/145000 = 50 per labor hour

Fabrication overhead rate = 7830000/290000 = 27 per machine hour

Overhead applied to Job Bravo = (50*22+22*27) = 1694